A super flying squirrel with a pink cape can fly 340 mph in still air can fly 200 miles downwind in the same amount of time that it can fly 140 miles upwind. The velocity of the wind is 60 mph. True or false?
Answer by mananth(12270)

distance downwind = 200 miles
speed downwind = 340+60
Time taken = 200/400 = 1/2 hour
.
.
distance traveled upwind =140 miles
speed = 340-60 = 280
time taken = 140/280 = 1/2 hour
Both the times are equal.
So it is true